# Susan Athey    Emil Palikot

### v4

@@ Appendix F Proportional hazards assumption
Note: Diagnostics of the proportional hazard model. The first three columns show results for the Mentoring model, and the last three for Challenges. City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is also a discrete variable with levels at: no education mentioned, non-traditional credential, college degree, masters, and postgraduate. Prof experience is the number of years between the start date of the first employment of program participants and 2023.
@@ Appendix G Bounds on the Average Treatment Effect Estimates
In this section, we perform robustness checks on the average treatment effect estimates. First, we consider several cases in which we change which subjects are included in the analysis. Second, we estimate AT bounds following lee2021bounding. Third, we estimate the average treatment effects at 4 months after the programs.
@@
We consider four specifications for Mentoring. First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where both candidates were assigned to treatment due to control participants securing early employment. Second, restricting to the 126 perfectly matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3). Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points. Fourth, additionally including the 9 non-participating mentors under the assumption their hypothetical mentees would not have found jobs reduces the estimate to 13.0 percentage points. All four specifications yield treatment effects between 11.9 and 16.4 percentage points, each statistically significant at the 5% level.
@@
   Balanced pairs only 126 126 0.603 0.484 0.119*
@@
Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications. We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of lee2009training. We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
The Lee bounding procedure addresses selective attrition by considering worst-case scenarios about which individuals are missing. Since treatment experienced higher attrition, the conservative assumption for estimating positive effects is that the best-performing treatment participants were most likely to attrite (e.g., successful job-finders who deactivated LinkedIn). To construct the lower bound, we trim 6.8% of observations from the upper tail of the observed treatment outcome distribution, the proportion representing excess attrition relative to control, and compare the mean of this trimmed distribution to the full control group mean. The upper bound uses all observed treatment outcomes compared to all observed control outcomes. We estimate Lee bounds of [0.065, 0.114]. The upper bound of 0.114 is nearly identical to our main point estimate of 0.113, while the positive lower bound of 0.065 indicates that even under worst-case assumptions about selective attrition, the Challenges program effect remains economically meaningful. The 95% confidence interval on the lower bound is [-0.029, 0.159]; while this marginally includes zero due to sampling variability in the observed outcomes, the point estimate is bounded away from zero.
@@
Our main analysis uses employment outcomes measured twelve months post-program. However, we were unable to collect LinkedIn data for 17 Challenges treatment participants and 7 control participants at the twelve-month measurement due to profile URL changes or privacy setting modifications. A natural concern is whether selective attrition biases our estimated treatment effects. To address this concern, we re-estimate average treatment effects using employment outcomes measured four months after program completion. At this earlier measurement point, we successfully collected complete LinkedIn data for all randomized participants in both programs. Any attrition-induced bias is therefore eliminated at this time horizon.
@@
Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition. The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts card2018works. If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
Appendix H Policy estimation algorithms
@@
2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
3. In the test set, construct predicted treatment effects using predictions from model $\mu_{train}$. Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p)_{train}-\hat{Y}(O)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p)_{train}$ is the predicted outcome under program $p$ constructed using $\mu_{train}$ and $\hat{Y}(O)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months. Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/t\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
@@ 2. In the training dataset, estimate outcome model using cross-validation: $\mu_{train}$
$\max_{z_{ip}}\sum_{i=1}^{I}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i=1}^{I}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{i=p}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
@@
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. See Appendix LABEL:cross_fit_appendix for details of the cross-fitting procedure. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@ H.1 Assignment Models
Using the training set, we estimate predicted treatment effects for each program relative to non-participation. We fit two causal survival forests (athey2019generalized; cui2023estimating): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. These settings follow standard recommendations in the generalized random forests (grf) package (tibshirani2024grf).
@@
Appendix I Evaluation Models
@@
For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (athey2019generalized; cui2023estimating). Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
Each causal survival forest uses identical parameters to the assignment models: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. Model fitting follows the grf package implementation (tibshirani2024grf).
The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively. The forest automatically constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a survival forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates. These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
@@
After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual. We then aggregate across the 100 repetitions: following chernozhukov2017double, we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$. Variance estimates are similarly aggregated by taking the median across repetitions.
@@
I.1 Model Validation: Recovering Experimental Effects
@@
For the Mentoring program (Panel a), AIPW predictions center tightly around the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03). The distribution is approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE. The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
For the Challenges program (Panel b), AIPW predictions show similarly strong performance, with a mean deviation of $0.06$ months (SE = 0.04). While the distribution exhibits slightly more dispersion than Mentoring—consistent with Challenges’ larger experimental sample and potentially greater heterogeneity—the mean deviation remains statistically indistinguishable from zero. The slight positive mean suggests modest underestimation, but the bias is economically negligible (less than 2% of the experimental ATE).

### v5

@@ Appendix F Proportional hazards assumption
Note: Diagnostics of the proportional hazard model. The first three columns show results for the Mentoring model, and the last three for Challenges. City is a discrete variable with levels: Small Town, Large City, and Warsaw; degree is a discrete variable indicating the highest level of education with levels: no education mentioned, non-traditional credential, college degree, masters, and postgraduate. Prof experience is the number of years between the start date of the first employment of program participants and 2023.
@@ Appendix G Bounds on the Average Treatment Effect Estimates
In this section, we perform robustness checks on the average treatment effect estimates. First, we consider several cases in which we change which subjects are included in the analysis. Second, we estimate AT bounds following Lee and Weidner (2021). Third, we estimate the average treatment effects at 4 months after the programs.
@@
We consider four specifications for Mentoring. First, our baseline estimate (15.3 percentage points, SE 5.7) uses 151 treatment and 147 control participants, reflecting the 4 mentor pairs where the mentees who were initially assigned to participate found jobs before the program start and were replaced by the control group. Second, restricting to the 147 matched pairs yields a slightly smaller but still substantial effect (11.9 percentage points, SE 6.3). Third, including the 4 participants who found jobs before program launch—coding their outcomes as 1—increases the estimated effect to 16.4 percentage points. Fourth, additionally including the 9 non-participating mentors under the assumption their hypothetical mentees would not have found jobs reduces the estimate to 13.0 percentage points. All four specifications yield treatment effects between 11.9 and 16.4 percentage points, each statistically significant at the 5% level.
@@
   Matched pairs only 147 147 0.603 0.484 0.119*
@@
Our twelve-month outcome measurement experienced modest attrition in the Challenges program: 6% had missing LinkedIn data due to profile URL changes or privacy setting modifications. We assess whether selective attrition could bias our estimated treatment effect using the bounding procedure of Lee (2009). We consider the most conservative case, in which subjects that found jobs before the start of the program are not considered; thus, we focus only on the attrition during the program, due to missing LinkedIn data.
The Lee bounding procedure addresses selective attrition by considering worst-case scenarios about which individuals are missing. Since treatment experienced higher attrition, the conservative assumption for estimating positive effects is that the best-performing treatment participants were most likely to attrite (e.g., successful job-finders who deactivated LinkedIn). To construct the lower bound, we trim 6.8% of observations from the upper tail of the observed treatment outcome distribution, the proportion representing excess attrition relative to control, and compare the mean of this trimmed distribution to the full control group mean. The upper bound uses all observed treatment outcomes compared to all observed control outcomes. Trimmed mean is calculated over the remaining observations.
We estimate Lee bounds of [0.065, 0.114]. The upper bound of 0.114 is nearly identical to our main point estimate of 0.113, while the positive lower bound of 0.065 indicates that even under worst-case assumptions about selective attrition, the Challenges program effect remains economically meaningful. The 95% confidence interval on the lower bound is [-0.029, 0.159]; while this marginally includes zero due to sampling variability in the observed outcomes, the point estimate is bounded away from zero.
@@
Our main analysis uses employment outcomes measured twelve months post-program. However, we were unable to collect LinkedIn data for 17 Challenges treatment participants and 7 control participants at the twelve-month measurement due to profile URL changes or privacy setting modifications. A natural concern is whether selective attrition biases our estimated treatment effects. To address this concern, we re-estimate average treatment effects using employment outcomes measured four months after program completion. At this earlier measurement point, we successfully collected complete LinkedIn data for all randomized participants in both programs. Any attrition-induced bias is therefore eliminated at this time horizon. Furthermore, comparing the estimates at 4 months with the main results shows persistence of the estimates over time.
@@
Importantly, these four-month estimates use the full experimental sample without any missing observations, confirming that our main findings are not an artifact of selective attrition. The consistency between four-month and twelve-month estimates also provides evidence that treatment effects do not attenuate over time—a key finding given that some labor market training programs show fadeout of impacts Card et al. (2018). If anything, the point estimates suggest persistent or growing effects, particularly for Challenges, though the conclusions are limited to 12 months from the start of each program.
Appendix H Stratified Estimation for Challenges
The Challenges experiment employed stratified randomization on track specialization (UX Design vs. Front-end Development), age category (over/under 30), and residence type (Warsaw, other large city, or small town). To account for this design, we re-estimate treatment effects using stratum-weighted estimators that properly reflect the randomization procedure.
Table 14 reports these stratum-reweighted estimates. For difference-in-means, we estimate treatment effects within each stratum and aggregate using stratum-size weights. For Cox proportional hazard models, we allow separate baseline hazards. Point estimates are similar to those in Table 2, with modestly smaller standard errors reflecting the efficiency gains from accounting for stratification.
Table 14: Average Treatment Effects on Employment Outcomes (Stratum-Reweighted)
Challenges (Stratum-Reweighted)
Difference-in-means Cox PH
New job Tech job Non-tech job Tech job
ATE 0.080 0.120** -0.040 0.111**
(0.051) (0.048) (0.042) (0.049)
ATE / Control mean (%) 16.0 45.2** -17.2 41.8**
(10.2) (18.2) (17.8) (18.4)
Treated 166 166 166 2,656
Control 218 218 218 3,488
Notes: Average treatment effects on employment outcomes using stratum-reweighted estimators. Columns 1–3 report stratum-weighted difference-in-means estimates; column 4 reports stratified Cox proportional hazard estimates. Standard errors in parentheses. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.
Appendix I Average Treatment Effect Estimates Using Survival Forest
This section estimates the average treatment effects of Mentoring and Challenges, on employment outcomes using causal survival forests. We use experimental samples only. The outcome of interest is the time until participants secured a new job, measured over a 13-month follow-up period. This approach estimates the difference in survival probabilities between treated and control groups at a specified time horizon. For each program, we fit a causal survival forest with 2,000 trees and compute the average treatment effect at 12 months post-intervention. Results are presented in Table 15.
Table 15: Average Treatment Effects of Mentoring and Challenges at 12 Months.
Mentoring Challenges
   ATE -5.180 -2.895
   SE (0.291) (0.244)
   N Treatment 151 183
   N Control 147 225
Note: Estimates of the average treatment effect of Mentoring and Challenges at 12 months. Estimates obtained using causal survival forests. Estimates based on experimental samples with no control variables.
The results indicate that the Mentoring program reduce the time it takes to find a job by 5.18 months,, while the Challenges program reduced this time by 2.9 months. Both estimates are statistically significant.
Figure 15 displays the average treatment effects of the Mentoring and Challenges programs on employment outcomes over time, measured from 3 to 13 months after program start. Both programs show negative treatment effects that grow larger in magnitude over time, indicating that participants in the treatment groups experience accelerated job finding compared to control groups. At 3 months, both programs show treatment effects near zero, suggesting minimal immediate impact. However, as time progresses, the effects become increasingly negative, meaning treated participants find jobs faster (have lower survival times). By month 13, the Mentoring program shows a treatment effect of approximately -6 months, while the Challenges program shows an effect of about -3.5 months.
Figure 15: Treatment Effect Estimates from Survival Forest.
Note: Note: Figure shows average treatment effects on employment probability at different time horizons following program completion. Treatment effects estimated using causal survival forests with 2,000 trees. Shaded regions represent 95% confidence intervals. Horizontal line at zero indicates no treatment effect.
Appendix J Policy estimation algorithms
@@ Appendix I Average Treatment Effect Estimates Using Survival Forest
2. In the training dataset, estimate outcome model $\mu_{train}$ and propensity model $e_{train}$.
3. In the test set, construct predicted treatment effects using predictions from models $\mu_{train}$ and $e_{train}$. Obtain $\tau^{Op,train}_{i,t}=E[\hat{Y}(p=P)_{train}-\hat{Y}(p=OutofDareIT)_{train}|X=x_{i,t},i\in I^{test}]$, where $\hat{Y}(p=P)_{train}$ is the predicted outcome under program $p\in$ (Mentoring, Challenges, Out of Dare IT) and $\hat{Y}(p=OutofDareIT)_{train}$ is the predicted outcome under Out of Dare IT and $t\in\left\{1,...,15\right\}$ denotes months. Compute mean treatment effects per user as: $\tau^{Op,train}_{i}=1/T\times\sum_{t=1}^{15}\tau^{Op,train}_{i,t}$,
@@
$\max_{z_{ip}}\sum_{i\in I^{test}}\sum_{p=1}^{P}z_{ip}\tau_{i}^{Op,train}\text{ s.t. }\sum_{i\in I^{test}}z_{ip}\leq Q^{p}\forall_{p}\text{ \& }\sum_{p=1}^{P}z_{ip}=1\forall_{i\in I^{test}}.$
@@
5. Using the test set, estimate new outcome and propensity models using cross-fitting and obtain predictions: $\hat{\mu}_{i,k}$ and $\hat{e}_{i,k}$ for all $i\in I^{test}$. Obtain $\hat{Y}_{i,k}(a^{*})$ the AIPW estimates of the predicted outcomes using cross-fitted models trained in the test set,
@@
Using the training set, we estimate predicted treatment effects for each program relative to non-participation. We fit two causal survival forests (Athey et al., 2019; Cui et al., 2023): one comparing Mentoring participants to applicants who did not participate in any program (including the control group of the Mentoring experiment), and another comparing Challenges participants to non-participants.
Each causal survival forest is estimated with the following parameters: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. Covariates include: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. These settings follow standard recommendations in the generalized random forests (grf) package (Tibshirani et al., 2024).
@@
Appendix K Evaluation Models
@@
For each policy evaluation, we estimate individual-level treatment effect predictions using causal survival forests (Athey et al., 2019; Cui et al., 2023). Within each cross-fitting fold (detailed below), we fit two causal survival forests: one comparing Mentoring participants to applicants who did not participate in any program, and another comparing Challenges participants to non-participants.
Each causal survival forest uses identical parameters to the assignment models: 2,000 trees, honest splitting, sample fraction of 0.5 per tree, minimum node size of 5, and horizon of 15 months. The covariate set is identical to that used in assignment models: STEM degree, residence type (Warsaw, small town), master’s degree, age over 30, professional experience (continuous), and UX specialization. Model fitting follows the grf package implementation (Tibshirani et al., 2024).
The causal survival forest produces augmented inverse propensity weighted (AIPW) scores $\psi_{i}^{M}$ and $\psi_{i}^{C}$ for each test-set individual, representing individual-level treatment effect estimates from Mentoring and Challenges relative to non-participation, respectively. The forest constructs these doubly-robust scores by combining: (1) propensity score estimates $\hat{e}(X_{i})$ from a regression probability forest predicting treatment assignment, and (2) conditional mean predictions $\hat{\mu}(X_{i},W)$ from honest regression forests predicting restricted mean survival time conditional on treatment status and covariates. These components are combined via the AIPW formula, yielding estimates that remain consistent if either the propensity or outcome model is correctly specified.
@@
After completing all folds in repetition $m$, each individual has predicted treatment effects $\psi_{i}^{M,m}$ and $\psi_{i}^{C,m}$ estimated on data that did not include that individual. We then aggregate across the 100 repetitions: we take the median prediction across repetitions for each individual: $\psi_{i}^{M}=\text{Median}(\psi_{i}^{M,1},\ldots,\psi_{i}^{M,100})$ and $\psi_{i}^{C}=\text{Median}(\psi_{i}^{C,1},\ldots,\psi_{i}^{C,100})$. Variance estimates are similarly aggregated by taking the median across repetitions. We follow the method described in Chernozhukov et al. (2017).
@@
K.1 Model Validation: Recovering Experimental Effects
@@
For the Mentoring program (Panel a), AIPW predictions center slightly above the RCT benchmark (SE, with a mean deviation of $0.08$ months (SE = 0.03). The difference between RCT and the mean deviation is statistically insignificant. The distribution appears approximately symmetric and concentrated near zero, indicating our cross-fitted estimates successfully recover the experimental ATE. The modest spread in the distribution reflects both the treatment effect heterogeneity among participants and sampling variation in the predictions.
For the Challenges program (Panel b), AIPW predictions show similarly good performance, with a mean deviation of $0.06$ months (SE = 0.04). While the distribution exhibits slightly more dispersion than Mentoring, the mean deviation remains statistically indistinguishable from the RCT mean.
@@
Appendix L Alternative Policy Evaluation Using Cross-Fitting
Our main off-policy evaluation in Section 6 employs a sample-splitting approach in which policy estimation and evaluation occur in separate datasets. This appendix presents an alternative approach using repeated cross-fitting following Fava (2025).
The procedure proceeds as follows. For each repetition $m\in\{1,\ldots,M\}$ with $M=100$, we randomly partition the full sample of applicants into $K=5$ folds, stratified by experimental group, using a different random seed. Within each repetition, we use $K$-fold cross-fitting: for each fold $k\in\{1,\ldots,K\}$, we train causal survival forests on the remaining K1K-1 folds, obtain out-of-sample predicted treatment effects $\hat{\tau}^{M}_{i}$ and $\hat{\tau}^{C}_{i}$ for observations in fold $k$, solve the constrained optimization problem in Equation 3 to determine optimal assignments, and compute AIPW scores under these assignments. After completing all folds within a repetition, we aggregate out-of-fold AIPW scores to obtain policy value estimates and standard errors computed as the sample standard deviation of individual scores divided by $\sqrt{n}$. Following Fava (2025) and Chernozhukov et al. (2018), we aggregate across the $M$ repetitions by taking the median of point estimates and the median of confidence interval bounds.
This approach requires a uniform convergence condition on the estimated treatment effects:
$\sup_{x\in\mathcal{X}}\left|\hat{\tau}^{p}(x)-\tau^{p}(x)\right|=o_{p}(1)\quad\text{for }p\in\{M,C\},$ (4)
where $\hat{\tau}^{p}(x)$ denotes the estimated conditional average treatment effect for program $p$ at covariate value $x$ and $\tau^{p}(x)$ is the true effect. This assumption is stronger than the pointwise consistency required by the sample-splitting approach in our main specification but enables valid inference when the same data are used for both policy learning and evaluation.
Table 16: Treatment Effects Under Targeted Allocation Policies: Repeated Cross-Fitting
Mentoring Challenges Overall
Panel A: Benchmark
Random allocation -2.22 (0.05) -2.54 (0.02) -2.38 (0.03)
Panel B: Targeted allocation policies
15% / 15% / 70% -5.26 (0.02) -3.94 (0.03) -4.60 (0.04)
33% / 33% / 34% -4.40 (0.05) -3.07 (0.04) -3.73 (0.04)
50% / 50% / 0% -3.41 (0.06) -2.55 (0.04) -2.98 (0.04)
Notes: Treatment effects measured as change in restricted mean survival time (months); negative values indicate faster job finding. Capacity allocations specify the share of applicants assigned to Mentoring/Challenges/Out. Random allocation assigns each applicant to either program with equal probability. Results are median estimates across 100 repetitions of cross-fitting with different random fold assignments. Standard errors in parentheses.
Table 16 presents median treatment effects across the 100 repetitions. Point estimates are similar to those in our main specification (Table 6): the overall treatment effect under the most selective policy (15/15/70) is 4.60-4.60 months compared to 4.34-4.34 months in the main analysis, while effects under expanded capacity (50/50/0) are 2.98-2.98 versus 2.86-2.86 months. The qualitative conclusions are unchanged: targeted allocation substantially outperforms random assignment, with gains from targeting diminishing as capacity expands.
The principal difference between specifications lies in the estimated standard errors. Cross-fitting yields smaller standard errors, particularly for selective policies where few applicants are assigned to each program. At 15% capacity, for example, the standard error on the Mentoring treatment effect is 0.02 compared to 0.09 in the main specification. This difference reflects efficiency gains from using the full sample for both estimation and evaluation rather than reserving half for each task. Additionally, the median aggregation across 100 repetitions improves reproducibility: different researchers using the same data but different random seeds will obtain identical results. These smaller standard errors, however, rely on the uniform convergence assumption holding in finite samples; the sample-splitting approach provides more conservative inference under weaker conditions. The concordance of point estimates across both approaches strengthens confidence in our conclusions regarding the value of targeted program assignment.